OpenStudy (anonymous):

Solve this system of equations {4y+3x=0 {y+x=1

OpenStudy (anonymous):

You need to multiply the 2nd equation by something that will either get rid of the y's or the x's... I will get rid of the y's so we need a -4y on the bottom 4y + 3x = 0 -4(y + x = 1) 4y + 3x = 0 -4y - 4x = -4 Now add the equations. -x = -4 x = 4 plug into one of the equations. y + x = 1 y + 4 = 1 y = -3 (4,-3) is the answer

OpenStudy (anonymous):

Thanks :)